A cylindrical can is to be constructed to hold a fixed volume of liquid. The cost of the


material used for the top and bottom of the can is 3 cents per square inch, and the cost of the


material used for the curved side is 2 cents per square inch. Use calculus to derive a simple


relationship between the radius and height of the can that is the least costly to construct.

1
Expert's answer
2021-11-25T16:58:11-0500

Let r be radius and h, height of the container.

Then;

"V = \pi ×r ^ 2 ×h"

and

"S(r) =(\pi ×r ^ 2) + (2\pi × r × h)"


Then the cost is;

"C(r) = 0.03\u03c0r\u00b2 +0.04 rh = C"

hence;

"h = (C - 0.03\pi ×r ^ 2)/(0.04\pi × r)"

and;

"V = (r(C - 0.03\pi × r ^ 2))/0.04"


The goal is to maximize V, since

"V= 25(C \u2013 0.09\u00b2)"


Then,

"V' = 0 \\

 r = (C/0.09π)^{1/2}\\

V" (r) = -4.5\u03c0r"


Hence,

V has maximum at "r =(\dfrac{C}{0.09 ×\pi })^ {1/2}"


Since;

"h/r = \dfrac{(C - 0.03\pi r ^ 2)}{(0.04\pi r ^ 2)}\\

r = 2/3h"